You deposited $20,000 in a savings account that pays 12 percent interest, compounded quarterly, planning to use it to finish your last year in college. Eighteen months later, you decide to go to the Rocky Mountains to become a ski instructor rather than continue in school, so you close out your account. How much money will you receive?
Answer #1

Number of quarter of period = 18/4 = 4.5
PV = 20000
Rate Per Quarter = 12%/4 = 3%

FV =PV*(1+r)^n = 20000*(1+12%/4)^4.5 = 22845.33